# CR3DT: Camera-RADAR Fusion for 3D Detection and Tracking

### v1

@@ Abstract
Accurate detection and tracking of surrounding objects is essential to enable self-driving vehicles. While Light Detection and Ranging (LiDAR) sensors have set the benchmark for high performance, the appeal of camera-only solutions lies in their cost-effectiveness. Notably, despite the prevalent use of Radio Detection and Ranging (RADAR) sensors in automotive systems, their potential in 3D detection and tracking has been largely disregarded due to data sparsity and measurement noise. As a recent development, the combination of RADARs and cameras is emerging as a promising solution. This paper presents Camera-RADAR 3D Detection and Tracking (CR3DT), a camera-RADAR fusion model for 3D object detection, and Multi-Object Tracking (MOT). Building upon the foundations of the State-of-the-Art (SotA) camera-only BEVDet architecture, CR3DT demonstrates substantial improvements in both detection and tracking capabilities, by incorporating the spatial and velocity information of the RADAR sensor. Experimental results demonstrate an absolute improvement in detection performance of 5.3% in mean Average Precision (mAP) and a 14.9% increase in Average Multi-Object Tracking Accuracy (AMOTA) on the nuScenes dataset when leveraging both modalities. CR3DT bridges the gap between high-performance and cost-effective perception systems in autonomous driving, by capitalizing on the ubiquitous presence of RADAR in automotive applications.
@@ I Introduction
Camera-only methods for cost-effectiveness: Substituting the expensive but highly performant LiDAR sensor with multiple cameras not only significantly reduces cost but might allow for wider adoption of self-driving technology within the automotive industry [9, 3] while delivering competitive performance.
@@
While both strategies have their merits, a distinct performance gap exists between camera-only and LiDAR-based models. Notably, in 3D detection tasks, SotA camera-only models achieve an mAP of 62.4 % [10], whilst SotA LiDAR-only models reach 69.5 % [11]. Similarly, SotA camera-only methods for 3D tracking achieve 65.3% AMOTA [12], whilst LiDAR-based methods score 71.5% AMOTA [13]. It is worth noting that these impressive performance numbers are reached by leveraging high-resolution image inputs, incorporating temporal information, or utilizing offline detections, as indicated in Table I. The latter implies that the model utilizes future data, rendering such detection systems impractical for real-time, i.e., online applications. These techniques may be used in various combinations, making fair comparisons about model performance hard to gauge (see Table I). This work targets the challenging task of object perception and tracking in autonomous driving and hence uses a computationally feasible, and fully online model that does not use additional temporal information (i.e., previous frames), to which we refer as the restricted model class, further detailed in Section IV.
@@
The RADAR sensor, prevalent in the automotive industry, has only recently emerged as a promising modality to bridge the existing performance gap. Previously, the sensor readings were deemed too noisy and sparse to improve 3D detection and tracking tasks [2, 15]. The absence of RADAR-based solutions in the nuScenes tracking challenge [16], a dataset that provides RADAR data, highlights this fact. While the sensor readings offer a spatial point cloud akin to LiDARs, although sparser, they also deliver richer measurements, capturing data points that include velocity and RADAR-reflectivity information. Only recently, RADAR has been used to support tasks such as in 3D segmentation [17] and 3D detection [18, 19]. Furthermore, different works have highlighted RADARs increased robustness to adverse weather conditions [15, 20, 21], especially when compared to LiDARs and cameras.
@@
This paper aims to bridge the performance gap between LiDAR and camera-only methods, by leveraging sensor fusion of the camera and RADAR modality. Thus, we introduce CR3DT a 3D detection and tracking solution synthesizing camera and RADAR data, capitalizing on the added velocity data of the RADAR readings. We opted to fuse these two types of sensor data within the Bird’s-Eye View (BEV) space, based on promising results demonstrated by using RADAR in this domain [17, 18, 19]. Additionally, the BEV space has yielded high-performance outcomes in both camera-only techniques [1, 9, 22, 23] and camera-LiDAR fusion methods [7, 24]. Furthermore, the BEV intermediate representation facilitates the incorporation of the 3D data from the RADAR, a notable advantage over other works that relied on 2D projections [21, 25]. To summarize the main points of this work:
@@
Detection performance evaluation: CR3DT achieves an mAP of 35.1 % and a nuScenes Detection Score (NDS) of 45.6 % on the nuScenes 3D detection validation set. This outperforms SotA single-frame, camera-only detection models within the restricted model class as defined in Section IV by 5.3 % mAP points. With the rich velocity information contained in RADAR data, the detector furthermore reduces the mean Average Velocity Error (mAVE) by 45.3 % versus the previously mentioned SotA camera-only detector.
@@ II Related Work
Popular autonomous driving datasets such as KITTI [4], Waymo [5], and nuScenes [2] allow model performance comparison of perception systems in the tasks of 3D object detection and tracking. Recent successful models typically leverage the BEV space to tackle these problems. In the following, we discuss previous LiDAR-based models, camera-only models, and camera-RADAR models addressing 3D object detection. Finally, we discuss the SotA models in 3D BEV space tracking.
@@ II-B Camera-Only 3D Object Detection
The search for 3D detection architectures has extensively focused on cost-effective camera-only models. When using camera-based systems in BEV space, 2D object detection techniques are being leveraged as much as possible due to their proven effectiveness. They are extended into the third dimension using multiple cameras spaced around the car. Recent development in this field yielded a technique called LSS [28], a pioneering method that leverages the camera intrinsics and extrinsic to project 2D image features into the BEV space of the car. While LSS utilizes the camera parameters as well as a per pixel attention-style Convolutional Neural Network (CNN) operation in BEV space, there exist other methods such as parameter-free lifting strategies [17] or methods based on deformable attention [9]. In this work, however, we restrict ourselves to the simpler LSS technique that is also utilized in the SotA BEVDet series.
Recent models leverage this image-view transformation to employ detection or segmentation heads on the aggregated BEV feature space [1, 9, 23]. They differ mostly in the way the image features are extracted and lifted, as well as subsequently encoded in the aggregated BEV space to be prepared for the detection or segmentation head. The CenterPoint architecture [8], while initially developed for the LiDAR modality, has been proven to work well with image-based features in BEV space, as shown by BEVDet [1], which denotes the current SotA in camera-only online 3D object detection within the restricted model class as defined in Section IV. Therefore, our proposed solution builds upon the BEVDet architecture, integrating the additional RADAR modality, and improving performance in terms of mAP, NDS, and mAVE.
@@ II-C Camera-RADAR Fusion Models
Recent camera-RADAR fusion architectures have demonstrated the potential of sensor-fusion within the 3D BEV space. SimpleBEV [17] represents an impactful work in 3D object segmentation and BEV-based sensor fusion, where the RADAR point clouds are rasterized and then concatenated with the BEV image-view features obtained from a ResNet-101 backbone. This concatenation fuses the spatial RADAR data with the feature-rich camera data after the projection of the image features into 3D BEV space. The proposed technique demonstrated an approximate 8%-point increase in absolute Intersection over Union (IOU) score for the nuScenes segmentation task. Following a similar idea, the BEVDet4D paradigm introduced in [23] shows the advantages of concatenating different sensor streams in BEV space, with the key difference that instead of RADAR, previous camera BEV features are fused with the current ones. Lastly, recent interest in cross-attention from the sensor-fusion community can be seen in [18], where deformable cross-attention can be found as a competitive method for BEV multi-modal fusion in the tasks of BEV segmentation and 3D detection and tracking. Hence, recent work strongly suggests that fusing RADAR with camera data yields a cost-effective performance increase across different perception tasks for autonomous driving. Following this trend, this work proposes a 3D detection and tracking architecture that evaluates different fusion strategies in BEV space and achieves significant improvement in detection and tracking performance metrics, when compared to the SotA baselines belonging to the same restricted model class.
@@ II-D Multi Object Tracking
In the context of 3D MOT, numerous solutions have traditionally relied on the LiDAR sensor modality due to its ability to provide a comprehensive 360-degree Field of View (FoV) [8, 29, 30, 31]. In contrast, 3D BEV tracking systems based on camera-only detection models have been relatively rare [32, 33]. This scarcity is, in part, because camera-based approaches often extend 2D MOT techniques, which rely on tracking by detection and utilize motion or appearance cues for association between frames. These 2D techniques inherently lack the capacity to leverage the rich spatial information provided by a 360-degree FoV sensor such as LiDAR.
More recent works in camera-only tracking have begun to diverge from 2D MOT techniques, focusing on associating temporal image-feature embeddings of the objects being tracked [14, 34]. These methods capitalize on quasi-dense similarity learning to enable robust and precise tracking. One notable development in this line of research is the CC-3DT [3] model, which innovatively extends tracking capabilities to handle joint associations between different cameras, thereby enabling tracking tasks to be performed through cross-camera correlation. We exploit the tracker from CC-3DT and enhance it, showing in Table IV that our tuning improves the out-of-the-box performance of the CC-3DT tracker across all considered metrics.
@@ III-A Sensor Fusion in BEV Space
Similar to PointPillars, the image features are lifted directly into 64 channel, single cell pillars in the BEV grid, leading to an image feature size of (64x128x128). The RADAR data is also directly aggregated (employing averaging) into single pillars, this time utilizing all 18 RADAR channels, including the point’s x, y, and z locations. This implicitly adds the centroid of the radar point cloud in each pillar to the list of features to be used. Notice that we did not augment the RADAR data in any way, in contrast to PointPillars augmentation efforts for LiDAR point clouds [6]. This is due to the fact that the RADAR point cloud already includes more information compared to LiDAR, as explored in [17]. Subsequently, the RADAR tensor of size (18x128x128) is concatenated to the image tensor of size (64x128x128) and both are fed into the BEV feature encoding layer as a ((64+18)x128x128) tensor. The ablation study discussed in Section V-B showed furthermore, that it is beneficial to add a residual connection to the output of the feature encoding layer of dimension (256x128x128), leading to a final input size to the CenterPoint [8] detection head of ((256+18)x128x128).
@@ III-B Tracker Architecture
Following the naming convention in CC-3DT [3], the detections are associated with the active tracks of the KF at time with a greedy assignment given an affinity matrix . The matrix is composed of the appearance embedding similarity matrix , the motion correlation matrix , and the location correlation matrix , weighted according to the following:
@@
where the centroid correlation and the state difference correlation are defined as in [3]. The new velocity correlation weighting is computed as:
@@
We refer to our extended implementation of CC-3DT with adapted thresholds, weights, and the introduction of our novel velocity similarity weight as CC-3DT++. The three changes are motivated by ablation studies, detailed in Table III(a), Table III(b), and Table III(c), respectively.
@@ IV Experimental Setup
This section details the setup and definitions of the detection and tracking baselines, as well as the training hardware that was utilized to train the described models. Within this work, both detection and tracking evaluations were performed on the well-known nuScenes dataset, as it contains both RADAR and camera data, necessary for our method. We tested our model on the nuScenes validation set, enabling comparison with related work. For computational reasons and to facilitate comparison, all models were trained without the usage of Class-Balanced Grouping and Sampling (CBGS).
@@ IV-B Training Hardware
The training was performed on a single GeForce RTX 3090. To replicate the training of the BEVDet network, which was performed on 8 Graphics Processing Units (GPUs) with a batch size of 8, training was performed with gradient accumulation over 8 steps. To alleviate the sparsity inherent to RADAR data, five sweeps were used, i.e., the sweep associated with the current camera frames as well as the four previous RADAR sweeps were accumulated. No additional temporal information was used in our fusion models.
@@ V-B Camera-Radar Fusion Ablation
To analyze the camera-RADAR fusion method in more detail, we compare different fusion architectures in Table II. Namely, we evaluate the performance of two different intermediate fusion approaches in Table I(a) and the benefit of our residual connection in Table I(b). Our intermediate fusion approaches were inspired by VoxelNet [26] and PointPillars [6], respectively. The first approach includes a voxelization of the lifted RGB features and pure RADAR sensor data into cubes of dimension , leading to an alternate feature size of ((64+18)x10x128x128) and consequently to the necessity of a bev_compressor module in the form of a 3D convolution, similar to the one employed in [26]. The latter approach is the one described in Section III, which forgoes the voxelization in the z dimension and the subsequent 3D convolution, and directly aggregates the lifted RGB features and pure RADAR sensor data into pillars, leading to the known feature size of ((64+18)x128x128). As can be seen in Table I(a), omitting the discretization in the z dimension and the additional bev_compressor in the model architecture leads to an increase in the models mAP and NDS as well as a decrease in its mAVE. Going forth with the better model, we add a residual connection to the output of our initial camera-RADAR fusion as seen in the system architecture overview in Fig. 1. Such a residual connection leads to a further increase in the mAP and NDS of our model, although at the cost of an increase in mAVE, as can be seen in Table I(b). It is worth noting though, that the mAVE still outclasses the camera-only BEVDet base model, as seen in Table I.

### v2

@@ Abstract
To enable self-driving vehicles accurate detection and tracking of surrounding objects is essential. While Light Detection and Ranging (LiDAR) sensors have set the benchmark for high-performance systems, the appeal of camera-only solutions lies in their cost-effectiveness. Notably, despite the prevalent use of Radio Detection and Ranging (RADAR) sensors in automotive systems, their potential in 3D detection and tracking has been largely disregarded due to data sparsity and measurement noise. As a recent development, the combination of RADARs and cameras is emerging as a promising solution. This paper presents Camera-RADAR 3D Detection and Tracking (CR3DT), a camera-RADAR fusion model for 3D object detection, and Multi-Object Tracking (MOT). Building upon the foundations of the State-of-the-Art (SotA) camera-only BEVDet architecture, CR3DT demonstrates substantial improvements in both detection and tracking capabilities, by incorporating the spatial and velocity information of the RADAR sensor. Experimental results demonstrate an absolute improvement in detection performance of 5.3% in mean Average Precision (mAP) and a 14.9% increase in Average Multi-Object Tracking Accuracy (AMOTA) on the nuScenes dataset when leveraging both modalities. CR3DT bridges the gap between high-performance and cost-effective perception systems in autonomous driving, by capitalizing on the ubiquitous presence of RADAR in automotive applications. The code is available at: https://github.com/ETH-PBL/CR3DT.
@@ I Introduction
Camera-only methods for cost-effectiveness: Substituting the expensive but highly performant LiDAR sensor with multiple cameras not only significantly reduces cost, but might allow for wider adoption of self-driving technology within the automotive industry [9, 3] while delivering competitive performance.
@@
While both strategies have their merits, a distinct performance gap exists between camera-only and LiDAR-based models. Notably, in 3D detection tasks, SotA camera-only models achieve an mAP of 62.4 % [10], whilst SotA LiDAR-only models reach 69.5 % [11]. Similarly, SotA camera-only methods for 3D tracking achieve 65.3% AMOTA [12], whilst LiDAR-based methods score 71.5% AMOTA [13]. It is worth noting that these impressive performance numbers are reached by leveraging high-resolution image inputs, incorporating temporal information, or utilizing offline detections, as indicated in Table I. The latter implies that the model utilizes future data, rendering such detection systems impractical for real-time, i.e., online applications. These techniques may be used in various combinations, making fair comparisons about model performance difficult (see Table I). This work targets the challenging task of object perception and tracking in autonomous driving and hence uses a computationally feasible, and fully online model that does not use additional temporal information (i.e., data corresponding to previous frames), to which we refer as the restricted model class, further detailed in Section IV.
@@
The RADAR sensor, prevalent in the automotive industry, has only recently emerged as a promising modality to bridge the existing performance gap. Previously, the sensor readings were deemed too noisy and sparse to improve 3D detection and tracking tasks [2, 15]. The absence of RADAR-based solutions in the nuScenes tracking challenge [16], a dataset that provides RADAR data, highlights this fact. While the sensor readings offer a spatial point cloud akin to LiDARs, although sparser, they also deliver richer measurements, capturing data points that include velocity and RADAR-reflectivity information. Only recently, RADAR has been used to support tasks such as 3D segmentation [17] and 3D detection [18, 19]. Furthermore, different works have highlighted the RADARs increased robustness to adverse weather conditions [15, 20, 21], especially when compared to LiDARs and cameras.
@@
This paper aims to bridge the performance gap between LiDAR and camera-only methods, by leveraging sensor fusion of the camera and RADAR modality. Thus, we introduce CR3DT a 3D detection and tracking solution synthesizing camera and RADAR data, capitalizing on the added velocity data of the RADAR readings. We opted to fuse these two types of sensor data within the Bird’s-Eye View (BEV) space, based on promising results demonstrated by using RADAR in this domain [17, 18, 19]. Additionally, the BEV space has yielded high-performance outcomes in both camera-only techniques [1, 9, 22, 23] and camera-LiDAR fusion methods [7, 24]. Furthermore, the BEV space as an intermediate representation facilitates the incorporation of the 3D data from the RADAR, a notable advantage over other works that relied on 2D projections [21, 25]. The main points of this work are summarized below:
@@
Detection performance evaluation: CR3DT achieves an mAP of 35.1 % and a nuScenes Detection Score (NDS) of 45.6 % on the nuScenes 3D detection validation set. This outperforms SotA single-frame, camera-only detection models within the restricted model class as defined in Section IV by 5.3 % mAP points. With the rich velocity information contained in the RADAR data, the detector furthermore reduces the mean Average Velocity Error (mAVE) by 45.3 % versus the previously mentioned SotA camera-only detector.
@@ II Related Work
Popular autonomous driving datasets such as KITTI [4], Waymo [5], and nuScenes [2] allow model performance comparison of perception systems in the tasks of 3D object detection and tracking. Recent successful models typically leverage the BEV space to tackle these problems. In the following, we discuss previous LiDAR-based models, camera-only models, and camera-RADAR models addressing 3D object detection. Finally, we discuss the SotA models operating in 3D BEV space to perform object tracking.
@@ II-B Camera-Only 3D Object Detection
The search for 3D detection architectures has extensively focused on cost-effective camera-only models. When using camera-based systems in BEV space, 2D object detection techniques are being leveraged as much as possible due to their proven effectiveness. They are extended into the third dimension using multiple cameras spaced around the car. Recent development in this field yielded a technique called LSS [28], a pioneering method that leverages the camera intrinsics and extrinsic to project 2D image features into the BEV space of the car. While LSS utilizes the camera parameters as well as a per pixel attention-style Convolutional Neural Network (CNN) operation in BEV space, there exist other methods such as parameter-free lifting [17] or methods based on deformable attention [9]. In this work, however, we restrict ourselves to the simpler LSS technique that is also utilized in the SotA BEVDet series.
Recent models leverage this image-view transformation to employ detection or segmentation heads on the aggregated BEV feature space [1, 9, 23]. They differ mostly in the way the image features are extracted and lifted, as well as subsequently encoded in the aggregated BEV space to be prepared for the detection or segmentation head. The CenterPoint architecture [8], while initially developed for the LiDAR modality, has been proven to work well with image-based features in BEV space, as shown by BEVDet [1], which denotes the current SotA in camera-only online 3D object detection within the restricted model class as defined in Section IV. Therefore, our proposed solution builds upon the BEVDet architecture, additionally integrating the RADAR modality, and improving performance in terms of mAP, NDS, and mAVE.
@@ II-C Camera-RADAR Fusion Models
Recent camera-RADAR fusion architectures have demonstrated the potential of sensor-fusion within the 3D BEV space. SimpleBEV [17] represents an impactful work in 3D object segmentation and BEV-based sensor fusion, where the RADAR point clouds are rasterized and then concatenated with the BEV image-view features obtained from a ResNet-101 backbone. This concatenation fuses the spatial RADAR data with the feature-rich camera data after the projection of the image features into 3D BEV space. The proposed technique demonstrated an approximate 8%-point increase in absolute Intersection over Union (IOU) score for the nuScenes segmentation task. Following a similar idea, the BEVDet4D paradigm introduced in [23] shows the advantages of concatenating different sensor streams in BEV space, with the key difference being, that instead of RADAR data previous camera BEV features are fused with the current ones. Lastly, recent interest in cross-attention from the sensor-fusion community can be seen in [18], where deformable cross-attention can be found as a competitive method for BEV space multi-modal fusion in the tasks of BEV segmentation and 3D detection and tracking. Hence, recent work strongly suggests that fusing RADAR with camera data yields a cost-effective performance increase across different perception tasks for autonomous driving. Following this trend, this work proposes a 3D detection and tracking architecture that evaluates different fusion strategies in BEV space and achieves significant improvement in detection and tracking performance metrics, when compared to the SotA baselines belonging to the same restricted model class.
@@ II-D Multi Object Tracking
In the context of 3D MOT, numerous solutions have traditionally relied on the LiDAR sensor modality due to its ability to provide a comprehensive 360-degree Field of View (FoV) [8, 29, 30, 31]. In contrast, 3D BEV tracking systems based on camera-only detection models have been relatively rare [32, 33]. This scarcity is, in part, because camera-based approaches often extend 2D MOT techniques, which rely on tracking by detection, and utilize motion or appearance cues for association between frames. These 2D techniques inherently lack the capacity to leverage the rich spatial information provided by a 360-degree FoV sensor such as LiDAR.
More recent works in camera-only tracking have begun to diverge from 2D MOT techniques, focusing on associating temporal image-feature embeddings of the objects being tracked [14, 34]. These methods capitalize on quasi-dense similarity learning to enable robust and precise tracking. One notable development in this line of research is the CC-3DT [3] model, which innovatively extends tracking capabilities to handle joint associations between different cameras, thereby enabling tracking tasks to be performed through cross-camera correlation. We exploit the tracker of CC-3DT and enhance it, demonstrating in Table IV that our tuning improves the out-of-the-box performance of the CC-3DT tracker across all considered metrics.
@@ III-A Sensor Fusion in BEV Space
Similar to PointPillars, the image features are lifted directly into 64-channel, single cell pillars in the BEV grid, leading to an image feature size of (64x128x128). The RADAR data is also directly aggregated (employing averaging) into single pillars, utilizing all 18 RADAR channels, including the point’s x, y, and z locations. This implicitly adds the centroid of the radar point cloud in each pillar to the list of features to be used. Notice that we did not augment the RADAR data in any way, in contrast to PointPillars augmentation efforts for LiDAR point clouds [6]. This is due to the fact that the RADAR point cloud already includes more information compared to LiDAR, as explored in [17]. Subsequently, the RADAR tensor of size (18x128x128) is concatenated to the image tensor of size (64x128x128) and both are fed into the BEV feature encoding layer as a ((64+18)x128x128) tensor. Furthermore, the ablation study discussed in Section V-B showed, that it is beneficial to add a residual connection to the output of the feature encoding layer of dimension (256x128x128), leading to a final input size to the CenterPoint [8] detection head of ((256+18)x128x128).
@@ III-B Tracker Architecture
Following the naming convention in CC-3DT [3], the detections are associated with the active tracks of the KF at time with a greedy assignment given an affinity matrix . The matrix is composed of the appearance embedding similarity matrix , the motion correlation matrix , and the location correlation matrix , weighted according to the following equation:
@@
where the centroid correlation and the state difference correlation are defined as in [3]. The new velocity correlation weight is computed as:
@@
We refer to our extended implementation of CC-3DT with adapted thresholds, weights, and the introduction of our novel velocity similarity weight as CC-3DT++. The three changes are motivated by ablation studies, detailed in Table IV(a), Table IV(b), and Table IV(c), respectively.
@@ IV Experimental Setup
This section details the setup and definitions of the detection and tracking baselines, as well as the training hardware that was utilized to train the described models. Within this work, both detection and tracking evaluations were performed on the well-known nuScenes dataset, as it contains both the RADAR and camera data, necessary for our method. We tested our model on the nuScenes validation set, enabling comparison with related work. For computational reasons and to facilitate comparison, all models were trained without the usage of Class-Balanced Grouping and Sampling (CBGS).
@@ IV-B Training Hardware
The training was performed on a single GeForce RTX 3090. To replicate the training of the BEVDet network, which was performed on 8 Graphics Processing Units (GPUs) with a batch size of 8, training was performed with gradient accumulation over 8 steps. To alleviate the sparsity inherent to RADAR data, five sweeps were used, i.e., the sweep associated with the current camera frames as well as the four previous RADAR sweeps were accumulated. Note that the RADAR sensor has a higher data rate than the camera, hence the radar sweeps that are being accumulated correspond to timestamps that are strictly after the timestamp of the previous image frame. No additional temporal information was used in our fusion models.
@@ V-B Camera-Radar Fusion Ablation
To analyze the camera-RADAR fusion method in more detail, we compare different fusion architectures in Table II. Namely, we evaluate the performance of two different intermediate fusion approaches in Table II(a) and the benefit of our residual connection in Table II(b). Our intermediate fusion approaches were inspired by VoxelNet [26] and PointPillars [6], respectively. The first approach includes a voxelization of the lifted RGB features and the pure RADAR sensor data into cubes of dimension , leading to an alternate feature size of ((64+18)x10x128x128) and consequently to the necessity of a bev_compressor module in the form of a 3D convolution, similar to the one employed in [26]. The latter approach is the one described in Section III, which forgoes the voxelization in the z dimension and the subsequent 3D convolution, and directly aggregates the lifted RGB features and pure RADAR sensor data into pillars, leading to the known feature size of ((64+18)x128x128). As can be seen in Table II(a), omitting the discretization in the z dimension and the additional bev_compressor in the model architecture leads to an increase in the models mAP and NDS as well as a decrease in its mAVE. Going further with the better model, we add a residual connection to the output of our initial camera-RADAR fusion as seen in the system architecture overview in Fig. 1. Such a residual connection leads to a further increase in the mAP and NDS of our model, although at the cost of an increase in mAVE, as can be seen in Table II(b). It is worth noting though, that the mAVE still outclasses the camera-only BEVDet base model, as seen in Table I.